OpenStudy (anonymous):

Integrate csc((πx)-1).

OpenStudy (anonymous):

\[\int\limits_{}\csc (\pi x-1)\]

ganeshie8 (ganeshie8):

As a start : substitute \(\rm u = \pi x-1\)

ganeshie8 (ganeshie8):

next see http://math2.org/math/integrals/more/csc.htm

OpenStudy (anonymous):

\[\frac{ -1}{ \pi } \ln \left| \csc (\pi x-1) +\cot (\pi x-1)\right|\]

OpenStudy (anonymous):

This is what I got. Is it correct? @ganeshie8

ganeshie8 (ganeshie8):

looks good!